The common stock and debt of Northern Sludge are valued at $58 million and $24 million, respectively. Investors currently require a 14% return on the common stock and an 6% return on the debt.

If Northern Sludge issues an additional $5 million of common stock and uses this money to retire debt, what is the new expected return on the stock? Assume that the change in capital structure does not affect the risk of the debt and that there are no taxes. (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

What is the expected return on the stock?
Answer #1

r(Assets) = [r(Debt) * w(Debt)] + [r(Equity) * w(Equity)]

= [0.06 * {24/(58+24)}] + [0.14 * {58/(58+24)}]

= 0.0176 + 0.0990 = 0.1166, or 11.66%

r(Equity) = r(Assets) + [(D/E) * {r(Assets) - r(Debt)}]

= 0.1166 + [{(24-5)/(58+5)} * (0.1166 - 0.06)]

= 0.1166 + [0.3016 * 0.0566]

= 0.1166 + 0.0171 = 0.1337, or 13.37%